Calculate the angular frequency of a mass on a spring which is oscillating back and forth with a period of 0.2 seconds per cycle
Answer #1

The period(T) and angular frequency(\omega) of a mass on a spring which is oscillating back and forth are related by

\omega=\frac{2\pi}{T}   

We are given T = 0.2 secs per cycle. We can find (\omega) using the above equation. Putting T = 0.2 secs/cycle

\omega=\frac{2\pi}{0.2} = 31.42\frac{rad}{sec}

Therefore the angular frequency(\omega) of the system \approx 31 rad/sec